OpenStudy (anonymous):

Write the quadratic function f(x) = 3x2 - 6x + 9 in vertex form.

OpenStudy (anonymous):

3(x - 1)^2 + 6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

3(x^2 - 2x + 3) = 3(x - 1)^2 + 2) = 3(x-1)^2 + 6 - i should have written the method as well

OpenStudy (anonymous):

haha yeah thank you that helped a lot!

OpenStudy (anonymous):

yw